What conic is described by the equation 8x^2 + 8y^2 - 4x + y = 0. please show steps. Thank you.
Answer by lwsshak3(11628)

What conic is described by the equation 8x^2 + 8y^2 - 4x + y = 0. please show steps.
..
8x^2 + 8y^2 - 4x + y = 0
completing the square
8(x^2-x/2+1/16)+8(y^2+y/8+1/256)=1/2+1/32=16/32+1/32=17/32
8(x-1/4)^2+8(y+1/16)^2=17/32
divide by 8
(x-1/4)^2+(y+1/16)^2=17/256
This equation is a circle with center at (1/4, -1/16) and radius=sqrt(17/256)
see graph below as a visual check on the answer.
..
y=(17/256-(x-1/4)^2)^.5-1/16